OpenStudy (anonymous):

Let f(x) = ln(1 + (1/x)). Show that f'(x) = (x^2 - 1)/ (x^3 +x).

OpenStudy (anonymous):

for starters \[1+\frac{1}{x}=\frac{x+1}{x}\] and \[\ln\left(\frac{x+1}{x}\right)=\ln(x+1)-\ln(x)\] therefore taking the derivative should be relatively painless

OpenStudy (anonymous):

I get the first part, but how does the ln( (x+1)/x) = ln(x+1)-ln(x) make sense?

OpenStudy (ranga):

It is one of the logarithmic identities: log(A/B) = log(A) - log(B) log(A * B) = log(A) + log(B) etc... (same if you replace log with ln)

OpenStudy (anonymous):

yes BUT that doesn't help me find the derivative at all. the derivative is (x^2 -1)/(x^3 -1). Using the identity leads me to a dead end.

OpenStudy (anonymous):

\[f(x)=\ln\left(\frac{x+1}{x}\right)=\ln(x+1)-\ln(x)\] \[f'(x)=\frac{1}{x+1}-\frac{1}{x}\] is a start

OpenStudy (anonymous):

ooh i think you wrote something different above

OpenStudy (anonymous):

\[\frac{x^2-1}{x^3-1}=\frac{(x+1)(x-1)}{(x-1)(x^2+x+1)}=\frac{x+1}{x^2+x+1}\]

OpenStudy (anonymous):

Actually I figured it out! Thanks for your help!